# Aligning Machine and Human Visual Representations across Abstraction Levels

### v3

@@ 4.1 AligNet
THINGS dataset objects. Similar to Muttenthaler et al. [2023a, b], we use the THINGS triplet odd-one-out dataset [Hebart et al., 2020] for learning an affine transformation into a (global) human object similarity space. The THINGS dataset can formaly be defined as $\mathcal{D}\coloneqq(\{a_{s},b_{s}\}\mid\{i_{s},j_{s},k_{s}\})_{s=1}^{n}$ which denotes a dataset of $n$ object triplets and corresponding human odd-one-out responses, where $\{a,b\}\subset\{i,j,k\}$ and $\{a,b\}$ is the object pair that was chosen by a human participant to have the highest similarity. Let ${\bm{X}}\in\mathbb{R}^{m\times p}$ be the teacher model representations for the $m=1854$ objects in the THINGS dataset. Note that each category in the THINGS dataset is represented by one object image. From ${\bm{X}}$ we can construct a similarity matrix for all object pairs ${\bm{S}}\coloneqq{\bm{X}}{\bm{X}}^{\top}\in\mathbb{R}^{m\times m}$, where $S_{ij}={\bm{x}}_{i}^{\top}{\bm{x}}_{j}$ is the representational similarity for objects $i,j$.
Linear transformation. Our goal is to learn an affine transformation into the THINGS human object similarity space of the form: ${\bm{x}}^{\prime}={\bm{W}}{\bm{x}}+{\bm{b}}$. Here, ${\bm{W}}\in\mathbb{R}^{p\times p}$ is a learned transformation matrix, ${\bm{b}}\in\mathbb{R}^{p}$ is a bias, and ${\bm{x}}\in\mathbb{R}^{p}$ is the neural network representation for a single object image in the THINGS dataset. We learn the affine transformation for the representation of the image encoder space of the SigLIP-So400m teacher model (see Sec. A.1.4 for details about the teacher model). Using this affine transformation, an entry in the pairwise similarity matrix ${\bm{S}}^{\prime}$—which represents the similarity between two object images $i,j$—can now be written as $S^{\prime}_{i,j}\coloneqq\left({\bm{W}}{\bm{x}}_{i}+{\bm{b}}\right)^{\top}\left({\bm{W}}{\bm{x}}_{j}+{\bm{b}}\right)$.
Uncertainty distillation. We mainly follow the optimization process introduced in Muttenthaler et al. [2023b]. However, we modify their approach by injecting uncertainty measures about human odd-one-out responses into the representation space of the teacher, using a recent approximate Bayesian inference method for learning object concepts from human behavior [Muttenthaler et al., 2022]. Thus, we replace the negative log-likelihood of the discrete human odd-one-out choices—which we refer to as hard-alignment—with the negative log-likelihood of the probabilities for the pairwise triplet similarities obtained from the Bayesian inference model—referred to as soft-alignment. Hard-alignment is defined as,
$odd-one-out prediction\displaystyle\mathcal{L}_{\mathrm{hard-align}}({\bm{S}}^{\prime})\coloneqq-\frac{1}{n}\sum_{s=1}^{n}\log\underbrace{q\left(\{a_{s},b_{s}\}|\{i_{s},j_{s},k_{s}\},{\bm{S}}^{\prime}\right)}_{\text{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{odd-one-out prediction}}},$ (1)
@@
Superordinate clusters. We embed the ImageNet train set in the transformed representation space of the teacher and cluster the representations into superordinate categories using $k$-Means clustering. We find the optimal number of clusters $k$ using the Elbow criterion. We use the clusters for generating triplets of distinct ImageNet images by always sampling two images from the same cluster and one image from a different cluster. For all triplets that we generate we identify their odd-one-out choice using the representations of the surrogate teacher model.
Human-like responses. The responses of the surrogate model simulate a dataset of human-like triplet odd-one-out responses. In addition to the discrete odd-one-out choices, the dataset includes the exact relationships among all pairwise similarities in a triplet obtained from the probability space of the teacher model. Thus, we now have access to soft choices.
@@
$\mathcal{L}_{\mathrm{alignet}}({\bm{S}}^{\prime},{\bm{S}}^{\dagger})\coloneqq\frac{1}{B}\sum_{s=1}^{B}\log\sigma\left(\left[S^{\prime}_{i,j},S^{\prime}_{i,j},S^{\prime}_{j,k}\right],\tau^{\prime}\right)_{s}-\sigma\left(\left[S^{\prime}_{i,j},S^{\prime}_{i,j},S^{\prime}_{j,k}\right],\tau^{\prime}\right)_{s}\log\sigma\left(\left[S^{\dagger}_{i,j},S^{\dagger}_{i,k},S^{\dagger}_{j,k}\right],\tau^{\dagger}\right)_{s},$
@@
where $\theta^{*}$ are the parameters of the pretrained student before fine-tuning and $\theta^{\dagger}$ are the parameters of the fine-tuned student. The right-hand side is similar to the $\ell_{2}$-regularization employed for learning the UD transformation. It tries to preserve the structure of the pretrained representation space as a function of $\lambda$ which determines the strength of the regularization.
@@ 4.2 Representational Similarity Analysis
Representational Similarity Analysis (RSA) is a well-established method for comparing neural network representations—extracted from an arbitrary layer of the model—to representations obtained from human behavior [Kriegeskorte et al., 2008a]. In RSA, one first obtains representational similarity matrices (RSMs) for the human behavioral judgments and for the neural network representations (more specific details can be found in the Supplementary Material). These RSMs measure the similarity between pairs of examples according to each source. As in previous work [Cichy et al., 2019; King et al., 2019; Hebart et al., 2020; Muttenthaler et al., 2023a, b], we flatten the upper triangular of human and model RSMs respectively and quantify their similarities using use the Spearman rank correlation coefficient. In contrast to the Pearson correlation coefficient, the Spearman rank correlation is scale-invariant and thus better suited to measure similarities of judgments obtained from different sources.
@@ 4.3 Levels
We collected a new multi-level similarity judgment dataset from $N=473$ human participants, which we named Levels. This dataset is based on the triplet odd-one-out task and included judgments on three different types of triplets: coarse-grained semantic, which require deciding on the odd-one-out in broadly different categories; fine-grained semantic, which involved discerning subtle distinctions within the same category; and class-boundary, which tested for the capacity to identify category boundaries. When predominantly one image (e.g., $i$) was selected to be the odd-one-out across multiple participants, it indicated that the other two images (e.g., $j$ and $k$), which were not selected, were closer to each other in the participants’ concept space than either was to the odd-one-out (see Sec. 4.3 and the Supplementary Material for details about the data collection). This dataset allowed us to evaluate whether neural network models respond similarly as humans for the same set of stimuli on various levels of abstraction, and to examine the extent to which the models capture the inherent uncertainty in human judgments, as inferred from response latencies.
@@ 4.4 Alignment with conceptual hierarchy
When analyzing alignment with the conceptual hierarchy, we use the original ImageNet category labels for the images [Deng et al., 2009]. ImageNet is structured according to the WordNet hierarchy, from which we extract basic- and superordinate-categories that align with the prior cognitive work.
@@
The synthetically created AligNet data will be made publicly available upon publication. We already made the Levels data available on GIN prior to publication: https://doi.org/10.12751/g-node.hg4tdz. Both the analysis and the experiment code for Levels are hosted on GitHub and archived on the Zenodo platform (see https://zenodo.org/doi/10.5281/zenodo.13749101).
@@ Acknowledgements
L.M., F.B., and K.R.M. were in part supported by the German Ministry for Education and Research (BMBF) under Grants 01IS14013A-E, 01GQ1115, 01GQ0850, 01IS18025A, 031L0207D, and 01IS18037A. K.R.M. was partly supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grants funded by the Korean government (MSIT) (No. 2019-0-00079, Artificial Intelligence Graduate School Program, Korea University and No. 2022-0-00984, Development of Artificial Intelligence Technology for Personalized Plug-and-Play Explanation and Verification of Explanation). B.S. and F.B. were supported by European Research Council Consolidator Grant ERC-2020-COG-101000972 and Deutsche Forschungsgemeinschaft (DFG) Grant 462752742. Correspondence should be addressed to L.M., K.R.M., and A.L. The authors would like to thank Aravindh Mahendran, Mehdi S. M. Sajjadi, Robert Geirhos, and Xiaohua Zhai for fruitful discussions and Ishita Dasgupta, Katherine Hermann, and Shakir Mohamed for helpful comments on earlier versions of the manuscript.
@@ A.1 Soft-alignment
This section is organized as follows: We start by describing how we transform model representations into a space that matches human similarity judgments about coarse-grained semantic object relations. We introduce a novel affine transformation that matches human similarity judgments and injects the uncertainties that humans assign to their triplet odd-one-out choices into a model’s representation space. Using these human-aligned representations, we sample triplets of ImageNet [Deng et al., 2009] images differently than uniform random sampling by clustering the representations into superordinate categories and using those clusters for data partitioning. Finally, after having created AligNet triplets, we can fine-tune models with a novel triplet loss object function.
@@ A.1.1 Representational alignment
Data. To increase the degree of alignment between human and neural network similarity spaces, we use the publicly available THINGS dataset, which is a large behavioral dataset of $4.7$ million unique triplet responses from $12{,}340$ human participants for $m=1854$ natural object images [Hebart et al., 2023] from the public THINGS object concept and image database [Hebart et al., 2019]. Let $\mathcal{D}_{\mathrm{things}}\coloneqq(\{a_{s},b_{s}\}\mid\{i_{s},j_{s},k_{s}\})_{s=1}^{n}$ be the dataset of object triplets and corresponding human responses, where $\{a,b\}\subset\{i,j,k\}$ and $\{a,b\}$ is the object pair that was chosen by a human participant to have the highest similarity.
Odd-one-out accuracy. The triplet odd-one-out task is a commonly used task in the cognitive sciences to measure human notions of object similarity [Fukuzawa et al., 1988; Robilotto and Zaidi, 2004; Hebart et al., 2020; Muttenthaler et al., 2022]. To measure the degree of alignment between human and neural network similarity judgments in the THINGS triplet task, we embed the $m=1854$ THINGS images into the representation space of a neural network with ${\bm{X}}\in\mathbb{R}^{m\times p}$. Given vector representations ${\bm{x}}_{1}$, ${\bm{x}}_{2}$, and ${\bm{x}}_{3}$ of the three images in a triplet, we first construct a similarity matrix ${\bm{S}}\in\mathbb{R}^{3\times 3}$ where $S_{i,j}\coloneqq{\bm{x}}_{i}^{\top}{\bm{x}}_{j}$ is the dot product between a pair of image representations. We identify the closest pair of images in the triplet as $\operatorname*{arg\,max}_{i,j>i}S_{i,j}$ with the remaining image being the odd-one-out. We define odd-one-out accuracy as the fraction of triplets where the odd-one-out “chosen by a model” is identical to the human odd-one-out choice. Thus, we want to find an affine transformation of the general form ${{\bm{x}}}^{\prime}={\bm{W}}{\bm{x}}+{\bm{b}}$ with ${{\bm{x}}}^{\prime}\in\mathbb{R}^{p}$ that maximizes this accuracy. The pairwise similarity can then be defined as $S^{\prime}_{ij}=\left({\bm{W}}{\bm{x}}_{i}+{\bm{b}}\right)^{\top}\left({\bm{W}}{\bm{x}}_{j}+{\bm{b}}\right)$.
@@
$odd-one-out prediction\displaystyle\mathcal{L}_{\mathrm{hard-align}}({\bm{S}}^{\prime})\coloneqq-\frac{1}{n}\sum_{s=1}^{n}\log\underbrace{q\left(\{a_{s},b_{s}\}|\{i_{s},j_{s},k_{s}\},{\bm{S}}^{\prime}\right)}_{\text{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{odd-one-out prediction}}}.$
Modelling human uncertainties. Since each triplet response is a discrete choice, we don’t have direct access to the uncertainties of a human participant over the objects in a triplet. Thus, the above loss function optimizes a transform to match the human choice but does not take into account the uncertainties over the three odd-one-out alternatives. However, it is possible to model these uncertainties using Variational Interpretable Concept Embeddings [VICE; Muttenthaler et al., 2022], a recently proposed, approximate Bayesian inference method for learning an interpretable object concept space from human similarity judgments. VICE has shown remarkable performance in predicting the (dis-)agreement in human similarity judgments for multiple similarity judgment datasets, including THINGS [Muttenthaler et al., 2022].
We train a VICE model on the official THINGS train triplet dataset using the (default) hyperparameters recommended by the authors. To capture the uncertainties in human triplet responses, VICE learns a mean, $\mu\in\mathbb{R}^{m\times d}$, and a variance, $\sigma\in\mathbb{R}^{m\times d}$, for each object image and each object dimension respectively. Therefore, the set of VICE parameters is defined as $\theta=\{\mu,\sigma\}$. VICE uses the reparameterization trick [Kingma and Welling, 2014; Rezende et al., 2014] to generate an embedding matrix ${\bm{Y}}\in\mathbb{R}^{m\times d}$, ${\bm{Y}}_{\theta,\epsilon}=\mu+\sigma\odot\epsilon$, where $\epsilon\in\mathbb{R}^{m\times d}$ is entrywise $\mathcal{N}(0,1)$, and $\odot$ denotes the Hadamard (element-wise) product.
After convergence, we can use a VICE model to obtain a posterior probability distribution for each triplet in the data. We approximate the probability distribution using a Monte Carlo estimate [Graves, 2011; Blundell et al., 2015; Blei et al., 2017] from $R$ samples ${\bm{Y}}^{(r)}={\bm{Y}}_{\hat{\theta},\epsilon^{(r)}}$ for $r=1,\ldots,R$, yielding
$Monte-Carlo estimate(s)\hat{p}\left(\{y_{s},z_{s}\}|\{i_{s},j_{s},k_{s}\}\right)\coloneqq\frac{1}{R}\sum_{r=1}^{R}\underbrace{p\left(\{y_{s},z_{s}\}|\{i_{s},j_{s},k_{s}\},{\bm{Y}}^{(r)}\right)}_{\text{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{Monte-Carlo estimate(s)}}},$
@@
Uncertainty distillation. Using the above soft alignment loss function instead of the vanilla hard alignment loss proposed by Muttenthaler et al. [2023a] for aligning the model representations with human global object similarity, the final objective function for the uncertainty distillation (UD) transformation is defined as follows,
@@
Although this expression is similar to the global transform defined in Muttenthaler et al. [2023b], we find it to yield equally strong downstream task performance as the gLocal transform proposed in Muttenthaler et al. [2023b] while predicting human uncertainties better than the global transform. It appears as if there is barely any trade-off between representational alignment and downstream task performance for using the UD, whereas Muttenthaler et al. [2023b] found that the global transform yields slightly better human alignment but worse downstream task performance compared to the gLocal transform. We use the UD transformation for generating human-like similarity judgments by transforming a model’s representation space with UD.

### v4

@@ 4.1 AligNet
THINGS dataset objects. We use the THINGS dataset [28] — a behavioral dataset of human responses in a triplet odd-one-out task for object images — to build a surrogate model (denoted as teacher) that approximates a human object similarity space. This similarity space is determined by the human odd-one-out responses from which various human concepts can be inferred [28, 54]. The THINGS dataset can formaly be defined as $\mathcal{D}\coloneqq(\{a_{s},b_{s}\}\mid\{i_{s},j_{s},k_{s}\})_{s=1}^{n}$ which denotes a dataset of $n$ object triplets and corresponding human odd-one-out responses, where $\{a,b\}\subset\{i,j,k\}$ and $\{a,b\}$ is the object pair that was chosen by a human participant to have the highest similarity. Let ${\bm{X}}\in\mathbb{R}^{m\times p}$ be the teacher model representations for the $m=1854$ objects in the THINGS dataset. Note that each category in the THINGS dataset is represented by one object image. From $\bm{X}$} we can construct a similarity matrix for all object pairs ${\bm{S}}\coloneqq{\bm{X}}{\bm{X}}^{\top}\in\mathbb{R}^{m\times m}$, where $S_{ij}={\bm{x}}_{i}^{\top}{\bm{x}}_{j}$ is the representational similarity for objects $i,j$.
Linear transformation. Our goal is to learn an affine transformation into the THINGS human object similarity space of the form: ${\bm{x}}^{\prime}={\bm{W}}{\bm{x}}+{\bm{b}}$. Here, ${\bm{W}}\in\mathbb{R}^{p\times p}$ is a learned transformation matrix, ${\bm{b}}\in\mathbb{R}^{p}$ is a bias, and ${\bm{x}}\in\mathbb{R}^{p}$ is the neural network representation for a single object image in the THINGS dataset. We learn the affine transformation for the representation of the image encoder space of the SigLIP-So400m teacher model (see the Methods for details about the teacher model). Using this affine transformation, an entry in the pairwise similarity matrix $\bm{S}$$^{\prime}$—which represents the similarity between two object images i,ji,j—can now be written as $S^{\prime}_{i,j}\coloneqq\left({\bm{W}}{\bm{x}}_{i}+{\bm{b}}\right)^{\top}\left({\bm{W}}{\bm{x}}_{j}+{\bm{b}}\right)$.
Uncertainty distillation. We mainly follow the optimization process introduced in Muttenthaler et al. [56]. However, we modify their approach by injecting uncertainty measures about human odd-one-out responses into the representation space of the teacher, using a recent approximate Bayesian inference method for learning object concepts from human behavior [54]. Thus, we replace the negative log-likelihood of the discrete human odd-one-out choices—which we refer to as hard-alignment—with the negative log-likelihood of the probabilities for the pairwise triplet similarities obtained from the Bayesian inference model—referred to as soft-alignment. Hard-alignment is defined as,
[…]
@@
Superordinate clusters. After we find a transformation that best matches the human responses, we embed the ImageNet train set in the transformed representation space of the teacher and cluster the representations into superordinate categories using $k$-Means clustering. We find the optimal number of clusters $k$ using the Elbow criterion. We use the clusters for generating triplets of distinct ImageNet images by always sampling two images from the same cluster and one image from a different cluster. For all triplets that we generate we identify their odd-one-out choice using the representations of the surrogate teacher model.
Human-aligned responses. The responses of the surrogate model simulate a dataset of human-aligned triplet odd-one-out responses. In addition to the discrete odd-one-out choices, the dataset includes the exact relationships among all pairwise similarities in a triplet obtained from the probability space of the teacher model. Thus, we now have access to soft choices.
@@
$\mathcal{L}_{\mathrm{alignet}}({\bm{S}}^{\prime},{\bm{S}}^{\dagger})\coloneqq\frac{1}{B}\sum_{s=1}^{B}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\sigma\left(\left[S^{\prime}_{i,j},S^{\prime}_{i,k},S^{\prime}_{j,k}\right],\tau^{\prime}\right)_{s}}}\log\sigma\left(\left[S^{\prime}_{i,j},S^{\prime}_{i,k},S^{\prime}_{j,k}\right],\tau^{\prime}\right)_{s}$-
$\sigma\left(\left[S^{\prime}_{i,j},S^{\prime}_{i,k},S^{\prime}_{j,k}\right],\tau^{\prime}\right)_{s}\log\sigma\left(\left[S^{\dagger}_{i,j},S^{\dagger}_{i,k},S^{\dagger}_{j,k}\right],\tau^{\dagger}\right)_{s},$
@@
where $\theta^{*}$ are the parameters of the pretrained student before fine-tuning and $\theta^{\dagger}$ are the parameters of the fine-tuned student. The right-hand side is similar to the $\ell_{2}$-regularization employed for learning the UD transformation. It tries to preserve the (fine-grained) structure of the pretrained representation space as a function of $\lambda$ which determines the strength of the regularization.
@@ 4.2 Representational Similarity Analysis
Representational Similarity Analysis (RSA) is a well-established method for comparing neural network representations—extracted from an arbitrary layer of the model—to representations obtained from human behavior [42]. In RSA, one first obtains representational similarity matrices (RSMs) for the human behavioral judgments and for the neural network representations (more specific details can be found in the SI). These RSMs measure the similarity between pairs of examples according to each source. As in previous work [14, 37, 28, 55, 56], we flatten the upper triangular of human and model RSMs respectively and quantify their similarities using use the Spearman rank correlation coefficient. In contrast to the Pearson correlation coefficient, the Spearman rank correlation is scale-invariant and thus better suited to measure similarities of judgments obtained from different sources.
@@ 4.3 Levels
We collected a new multi-level similarity judgment dataset from $N=473$ human participants, which we named Levels. The dataset contains odd-one-out judgments on three different types of triplets: coarse-grained semantic, which require deciding on the odd-one-out in broadly different categories; fine-grained semantic, which involved discerning subtle within category distinctions; and class-boundary, which tested for category boundary detection. Consistent selection of the same odd-one-out image (e.g., $i$) in multiple participants indicated that the remaining two images (e.g., $j$ and $k$) were closer to each other in the participants’ concept space than either was to the odd-one-out (see the SI for details about the data collection). Levels allowed us to evaluate model-human alignment for the same set of stimuli on various levels of abstraction, and to assess how well the models capture the inherent uncertainty in human judgments, inferred from response latencies.
Human-to-human alignment. We computed the human noise ceiling for each abstraction setting in Levels using a Leave-One-Out (LOO) cross-validation approach. In LOO, the agreement level for a triplet is computed as the average match rate between a held-out participant’s response and the majority response of the remaining population. Thus, for a triplet that was used in five participants, on each LOO iteration, one participant response is held out and the remaining four comprise the population. The human-to-human reliability score is then calculated as the average agreement level across all triplets in the dataset.
@@ 4.4 Alignment with conceptual hierarchy
When analyzing alignment with the conceptual hierarchy, we use the original ImageNet category labels for the images [16]. ImageNet is structured according to the WordNet hierarchy, from which we extract basic- and superordinate-categories that align with the prior cognitive work. We quantify these changes with mixed-effects linear regressions, see SI B.3.2.
Code Availability
The training code is available at https://github.com/google-deepmind/alignet. The aligned model checkpoints are publicly available at https://console.cloud.google.com/storage/browser/alignet. Both the experiment and analysis code for Levels and for the human validation of the representational differences between AligNet and UnaligNet are hosted on GitHub and archived on the Zenodo platform (see https://zenodo.org/records/15554034 and https://zenodo.org/records/15554174).
@@
The synthetically created AligNet data are publicly available at https://console.cloud.google.com/storage/browser/alignet. The Levels data are available on GIN prior to publication: https://doi.org/10.12751/g-node.hg4tdz.
@@ Acknowledgements
L.M., F.B., and K.R.M. were in part supported by the German Ministry for Education and Research (BMBF) under Grants 01IS14013A-E, 01GQ1115, 01GQ0850, 01IS18025A, 031L0207D, and 01IS18037A. K.R.M. was partly supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grants funded by the Korean government (MSIT) (No. 2019-0-00079, Artificial Intelligence Graduate School Program, Korea University and No. 2022-0-00984, Development of Artificial Intelligence Technology for Personalized Plug-and-Play Explanation and Verification of Explanation). B.S. and F.B. were supported by European Research Council Consolidator Grant ERC-2020-COG-101000972 and Deutsche Forschungsgemeinschaft (DFG) Grant 462752742. Correspondence should be addressed to L.M., K.R.M., and A.L. The authors would like to thank Aravindh Mahendran, Mehdi S. M. Sajjadi, Robert Geirhos, and Xiaohua Zhai for fruitful discussions and Ishita Dasgupta, Katherine Hermann, Shakir Mohamed and the editors and reviewers for helpful comments on earlier versions of the manuscript.
@@ A.1 Soft-alignment
This section is organized as follows: We start by describing how we transform model representations into a space that matches human similarity judgments about coarse-grained semantic object relations. We introduce a novel affine transformation that matches human similarity judgments and injects the uncertainties that humans assign to their triplet odd-one-out choices into a model’s representation space. Using these human-aligned representations, we sample triplets of ImageNet [16] images differently than uniform random sampling by clustering the representations into superordinate categories and using those clusters for data partitioning. Finally, after having created AligNet triplets, we can fine-tune models with a novel triplet loss object function.
@@ A.1.1 Representational alignment
Data. To increase the degree of alignment between human and neural network similarity spaces, we use the publicly available THINGS dataset, which is a large behavioral dataset of $4.7$ million unique triplet responses from 12,34012{,}340 human participants for $m=1854$ natural object images [29] from the public THINGS object concept and image database [27]. Let $\mathcal{D}_{\mathrm{things}}\coloneqq(\{a_{s},b_{s}\}\mid\{i_{s},j_{s},k_{s}\})_{s=1}^{n}$ be the dataset of object triplets and corresponding human responses, where $\{a,b\}\subset\{i,j,k\}$ and $\{a,b\}$ is the object pair that was chosen by a human participant to have the highest similarity.
Odd-one-out accuracy. The triplet odd-one-out task is a commonly used task in the cognitive sciences to measure human notions of object similarity [20, 71, 28, 54]. To measure the degree of alignment between human and neural network similarity judgments in the THINGS triplet task, we embed the $m=1854$ THINGS images into the representation space of a neural network with ${\bm{X}}\in\mathbb{R}^{m\times p}$. Given vector representations ${\bm{x}}_{1}$, ${\bm{x}}_{2}$, and ${\bm{x}}_{3}$ of the three images in a triplet, we first construct a similarity matrix ${\bm{S}}\in\mathbb{R}^{3\times 3}$ where $S_{i,j}\coloneqq{\bm{x}}_{i}^{\top}{\bm{x}}_{j}$ is the dot product between a pair of image representations. We identify the closest pair of images in the triplet as $\operatorname*{arg\,max}_{i,j>i}S_{i,j}$ with the remaining image being the odd-one-out. We define odd-one-out accuracy as the fraction of triplets where the odd-one-out “chosen by a model” is identical to the human odd-one-out choice. Thus, we want to find an affine transformation of the general form ${{\bm{x}}}^{\prime}={\bm{W}}{\bm{x}}+{\bm{b}}$ with ${{\bm{x}}}^{\prime}\in\mathbb{R}^{p}$ that maximizes this accuracy. The pairwise similarity can then be defined as $S^{\prime}_{ij}=\left({\bm{W}}{\bm{x}}_{i}+{\bm{b}}\right)^{\top}\left({\bm{W}}{\bm{x}}_{j}+{\bm{b}}\right)$.
@@
[…]
Modelling human uncertainties. Since each triplet response is a discrete choice, we don’t have direct access to the uncertainties of a human participant over the objects in a triplet. Thus, the above loss function optimizes a transform to match the human choice but does not take into account the uncertainties over the three odd-one-out alternatives. However, it is possible to model these uncertainties using Variational Interpretable Concept Embeddings [VICE; 54], a recently proposed, approximate Bayesian inference method for learning an interpretable object concept space from human similarity judgments. VICE has shown remarkable performance in predicting the (dis-)agreement in human similarity judgments for multiple similarity judgment datasets, including THINGS [54].
We train a VICE model on the official THINGS train triplet dataset using the (default) hyperparameters recommended by the authors. To capture the uncertainties in human triplet responses, VICE learns a mean, $\mu\in\mathbb{R}^{m\times d}$, and a variance, $\sigma\in\mathbb{R}^{m\times d}$, for each object image and each object dimension respectively. Therefore, the set of VICE parameters is defined as $\theta=\{\mu,\sigma\}$. VICE uses the reparameterization trick [38, 69] to generate an embedding matrix ${\bm{Y}}\in\mathbb{R}^{m\times d}$, ${\bm{Y}}_{\theta,\epsilon}=\mu+\sigma\odot\epsilon$, where $\epsilon\in\mathbb{R}^{m\times d}$ is entrywise $\mathcal{N}(0,1)$, and $\odot$ denotes the Hadamard (element-wise) product.
After convergence, we can use a VICE model to obtain a posterior probability distribution for each triplet in the data. We approximate the probability distribution using a Monte Carlo estimate [26, 8, 7] from $R$ samples ${\bm{Y}}^{(r)}={\bm{Y}}_{\hat{\theta},\epsilon^{(r)}}$ for $r=1,\ldots,R$, yielding
[…]
@@
Uncertainty distillation. Using the above soft alignment loss function instead of the vanilla hard alignment loss proposed by Muttenthaler et al. [55] for aligning the model representations with human global object similarity, the final objective function for the uncertainty distillation (UD) transformation is defined as follows,
@@
Although this expression is similar to the global transform defined in Muttenthaler et al. [56], we find it to yield equally strong downstream task performance as the gLocal transform proposed in Muttenthaler et al. [56] while predicting human uncertainties better than the global transform. It appears as if there is barely any trade-off between representational alignment and downstream task performance for using the UD, whereas Muttenthaler et al. [56] found that the global transform yields slightly better human alignment but worse downstream task performance compared to the gLocal transform. We use the UD transformation for generating human-like similarity judgments by transforming a model’s representation space with UD.